# Concept-Based Interpretable Reinforcement Learning with Limited to No Human Labels

### v1

@@ B.2 Budget Allocation Effectiveness
$B$ $M$ $R$uparrow $c$ Error $\downarrow$
PixelCartPole-v1 300 1 0.276 344±0.104 789plus-or-minus0.2763440.104789\$0.276\,344$\pm$\$0.104\,789\$0.276 344 ± 0.104 789 0.153 573±0.065 982plus-or-minus0.1535730.065982\$0.153\,573$\pm$\$0.065\,982\$0.153 573 ± 0.065 982
@@
DoorKey-7x7 100 1 0.765 391±0.039 967plus-or-minus0.7653910.039967\$0.765\,391$\pm$\$0.039\,967\$0.765 391 ± 0.039 967 0.260 404±0.041 564plus-or-minus0.2604040.041564\$0.260\,404$\pm$\$0.$0 4$ 0.041 564
@@
DynamicObstacles-5x5 100 1 0.957 439±0.020 369plus-or-minus0.9574390.020369\$0.957\,439$\pm$\$0.020\,369\$0.957 439 ± 0.020 369 0.040 794±0.006 774plus-or-minus0.0407940.006774\$0.040\,794$\pm$\$0.006\,774\$0.040 794 ± 0.006 774
@@
Table 7: Performance of LICORICE on all environments for varying budgets. We select $M$ through grid search, optimizing $R$ $c$ error without additional weighting since they are observed to have the same magnitude. The reward is reported as the fraction of the reward upper bound. For PixelCartPole-v1, $c$ error is MSE; for the other environments, $c$ error is 1 - accuracy. The $\pm$ [value] part shows the standard deviation. This shows a more complete version of the results in Table 2.
In Table 7, we present an extension of the results in Table 2, including the standard deviation. As expected, as the budget increases, the standard deviation for both the reward and concept error tends to decrease. The one exception is the reward for PixelCartPole-v1. Interestingly, the standard deviation is highest for $B=400$. We suspect this is because the concept errors may be more critical here, leading to higher variance in the reward performance.
@@
PixelCartPole-v1 300 1 0.064 712±0.017 953plus-or-minus0.0647120.017953\$0.064\,712$\pm$\$0.017\,953\$0.064 712 ± 0.017 953 0.170 068±0.188 407plus-or-minus0.1700680.188407\$0.170\,068$\pm$\$0.188\,407\$0.170 068 ± 0.188 407 0.193 75±0.248 575plus-or-minus0.193750.248575\$0.193\,75$\pm$\$0.248\,575\$0.193 75 ± 0.248 575
@@
DoorKey-7x7 100 1 0.711 403±0.042 15plus-or-minus0.7114030.04215\$0.711\,403$\pm$\$0.042\,15\$0.711 403 ± 0.042 15 0.450 326±0.036 168plus-or-minus0.4503260.036168\$0.450\,326$\pm$\$0.036\,168\$0.450 326 ± 0.036 168 0.313±0.034 549plus-or-minus0.3130.034549\$0.313$\pm$\$0.034\,549\$0.313 ± 0.034 549
@@
DynamicObstacles-5x5 100 1 0.268 711±0.373 027plus-or-minus0.2687110.373027\$0.268\,711$\pm$\$0.373\,027\$0.268 711 ± 0.373 027 0.200 711±0.0328plus-or-minus0.2007110.0328\$0.200\,711$\pm$\$0.0328\$0.200 711 ± 0.0328 0.159 636±0.029 329plus-or-minus0.1596360.029329\$0.159\,636$\pm$\$0.029\,329\$0.159 636 ± 0.029 329
@@
Table 8: Performance of LICORICE with GPT-4o integrated into the loop for all environments across different budgets, along with the concept labeling error of GPT-4o as a reference. We select $M$ through grid search, optimizing $R$ $c$ error without additional weighting since they are observed to have the same magnitude. GPT-4o $c$ error is evaluated on a random sample of 50 observations from the same rollout set used for LICORICE+GPT-4o, due to API cost constraints. This shows a more complete version of the results in Table 3.

### v2

@@ B.2 Budget Allocation Effectiveness
Environment $B$ $M$ $R$uparrow $c$ Error $\downarrow$
PixelCartPole 300 1 0.276 344±0.104 789plus-or-minus0.2763440.104789\$0.276\,344$\pm$\$0.104\,789\$0.276 344 ± 0.104 789 0.153 573±0.065 982plus-or-minus0.1535730.065982\$0.153\,573$\pm$\$0.065\,982\$0.153 573 ± 0.065 982
@@
DoorKey 100 1 0.765 391±0.039 967plus-or-minus0.7653910.039967\$0.765\,391$\pm$\$0.039\,967\$0.765 391 ± 0.039 967 0.260 404±0.041 564plus-or-minus0.2604040.041564\$0.260\,404$\pm$\$0.$0 4$ 0.041 564
@@
DynamicObstacles 100 1 0.957 439±0.020 369plus-or-minus0.9574390.020369\$0.957\,439$\pm$\$0.020\,369\$0.957 439 ± 0.020 369 0.040 794±0.006 774plus-or-minus0.0407940.006774\$0.040\,794$\pm$\$0.006\,774\$0.040 794 ± 0.006 774
@@
Boxing 1000 5 0.942 475±0.050 04plus-or-minus0.9424750.05004\$0.942\,475$\pm$\$0.050\,04\$0.942 475 ± 0.050 04 0.108 584±0.008 43plus-or-minus0.1085840.00843\$0.108\,584$\pm$\$0.008\,43\$0.108 584 ± 0.008 43
2000 5 0.980 631±0.047 948plus-or-minus0.9806310.047948\$0.980\,631$\pm$\$0.047\,948\$0.980 631 ± 0.047 948 0.098 675±0.059 967plus-or-minus0.0986750.059967\$0.098\,675$\pm$\$0.059\,967\$0.098 675 ± 0.059 967
3000 5 1.004 997±0.017 713plus-or-minus1.0049970.017713\$1.004\,997$\pm$\$0.017\,713\$1.004 997 ± 0.017 713 0.049±0.010 967plus-or-minus0.0490.010967\$0.049$\pm$\$0.010\,967\$0.049 ± 0.010 967
Pong 1000 5 0.994 762±0.002 988plus-or-minus0.9947620.002988\$0.994\,762$\pm$\$0.002\,988\$0.994 762 ± 0.002 988 0.647 428±0.041 13plus-or-minus0.6474280.04113\$0.647\,428$\pm$\$0.041\,13\$0.647 428 ± 0.041 13
3000 5 0.995 238±0.004 836plus-or-minus0.9952380.004836\$0.995\,238$\pm$\$0.004\,836\$0.995 238 ± 0.004 836 0.428 171±0.046 414plus-or-minus0.4281710.046414\$0.428\,171$\pm$\$0.046\,414\$0.428 171 ± 0.046 414
5000 5 0.996 333±0.018 424plus-or-minus0.9963330.018424\$0.996\,333$\pm$\$0.018\,424\$0.996 333 ± 0.018 424 0.243 221±0.162 857plus-or-minus0.2432210.162857\$0.243\,221$\pm$\$0.162\,857\$0.243 221 ± 0.162 857
Table 6: Performance of LICORICE on all environments for varying budgets. We select $M$ to optimize $R$ $c$ error without additional weighting since they are observed to have the same magnitude. The reward is reported as the fraction of the reward upper bound. For PixelCartPole, $c$ error is MSE; for the other environments, $c$ error is 1 - accuracy. The $\pm$ [value] part shows the standard deviation. This shows a more complete version of the results in Figure 3.
In Table 6, we present an extension of the results in Figure 3, including the standard deviation. As expected, as the budget increases, the standard deviation for both the reward and concept error tends to decrease. The one exception is the reward for PixelCartPole. Interestingly, the standard deviation is highest for $B=400$. We suspect this is because the concept errors may be more critical here, leading to higher variance in the reward performance.
@@
PixelCartPole 300 1 0.064 712±0.017 953plus-or-minus0.0647120.017953\$0.064\,712$\pm$\$0.017\,953\$0.064 712 ± 0.017 953 0.170 068±0.188 407plus-or-minus0.1700680.188407\$0.170\,068$\pm$\$0.188\,407\$0.170 068 ± 0.188 407 0.193 75±0.248 575plus-or-minus0.193750.248575\$0.193\,75$\pm$\$0.248\,575\$0.193 75 ± 0.248 575
@@
DoorKey 100 1 0.711 403±0.042 15plus-or-minus0.7114030.04215\$0.711\,403$\pm$\$0.042\,15\$0.711 403 ± 0.042 15 0.450 326±0.036 168plus-or-minus0.4503260.036168\$0.450\,326$\pm$\$0.036\,168\$0.450 326 ± 0.036 168 0.313±0.034 549plus-or-minus0.3130.034549\$0.313$\pm$\$0.034\,549\$0.313 ± 0.034 549
@@
DynamicObstacles 100 1 0.268 711±0.373 027plus-or-minus0.2687110.373027\$0.268\,711$\pm$\$0.373\,027\$0.268 711 ± 0.373 027 0.200 711±0.0328plus-or-minus0.2007110.0328\$0.200\,711$\pm$\$0.0328\$0.200 711 ± 0.0328 0.159 636±0.029 329plus-or-minus0.1596360.029329\$0.159\,636$\pm$\$0.029\,329\$0.159 636 ± 0.029 329
@@
Boxing 1000 5 0.342 437±0.088 819plus-or-minus0.3424370.088819\$0.342\,437$\pm$\$0.088\,819\$0.342 437 ± 0.088 819 0.978 089±0.011 367plus-or-minus0.9780890.011367\$0.978\,089$\pm$\$0.011\,367\$0.978 089 ± 0.011 367 0.781 24±0.005 29plus-or-minus0.781240.00529\$0.781\,24$\pm$\$0.005\,29\$0.781 24 ± 0.005 29
2000 5 0.228 937±0.107 898plus-or-minus0.2289370.107898\$0.228\,937$\pm$\$0.107\,898\$0.228 937 ± 0.107 898 0.984 122±0.006 578plus-or-minus0.9841220.006578\$0.984\,122$\pm$\$0.006\,578\$0.984 122 ± 0.006 578 0.794 52±0.014 46plus-or-minus0.794520.01446\$0.794\,52$\pm$\$0.014\,46\$0.794 52 ± 0.014 46
3000 5 0.221 577±0.009 677plus-or-minus0.2215770.009677\$0.221\,577$\pm$\$0.009\,677\$0.221 577 ± 0.009 677 0.976 101±0.009 001plus-or-minus0.9761010.009001\$0.976\,101$\pm$\$0.009\,001\$0.976 101 ± 0.009 001 0.790 14±0.016 773plus-or-minus0.790140.016773\$0.790\,14$\pm$\$0.016\,773\$0.790 14 ± 0.016 773
Pong 1000 5 0.004 167±0.002 525plus-or-minus0.0041670.002525\$0.004\,167$\pm$\$0.002\,525\$0.004 167 ± 0.002 525 0.794 917±0.083 28plus-or-minus0.7949170.08328\$0.794\,917$\pm$\$0.083\,28\$0.794 917 ± 0.083 28 0.883 29±0.005 45plus-or-minus0.883290.00545\$0.883\,29$\pm$\$0.005\,45\$0.883 29 ± 0.005 45
3000 5 $0.0$\pm$0.0$$ 0.737 935±0.070 374plus-or-minus0.7379350.070374\$0.737\,935$\pm$\$0.070\,374\$0.737 935 ± 0.070 374 0.860 15±0.031 72plus-or-minus0.860150.03172\$0.860\,15$\pm$\$0.031\,72\$0.860 15 ± 0.031 72
5000 5 0.126 19±0.170 042plus-or-minus0.126190.170042\$0.126\,19$\pm$\$0.170\,042\$0.126 19 ± 0.170 042 0.706 473±0.009 016plus-or-minus0.7064730.009016\$0.706\,473$\pm$\$0.009\,016\$0.706 473 ± 0.009 016 0.874 07±0.021 957plus-or-minus0.874070.021957\$0.874\,07$\pm$\$0.021\,957\$0.874 07 ± 0.021 957
@@
Table 7: Performance of LICORICE with GPT-4o integrated into the loop for all environments across different budgets, along with the concept labeling error of GPT-4o as a reference. We select $M$ to optimize $R$ $c$ error without additional weighting since they are observed to have the same magnitude. GPT-4o $c$ error is evaluated on a random sample of 50 observations from the same rollout set used for LICORICE+GPT-4o, due to API cost constraints. This shows a more complete version of the results in Figure 3.
Environment $B$ Sequential-Q Disagreement-Q Random-Q
$R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$
PixelCartPole 300 0.172 72±0.029 967plus-or-minus0.172720.029967\$0.172\,72$\pm$\$0.029\,967\$0.172 72 ± 0.029 967 0.113 868±0.032 088plus-or-minus0.1138680.032088\$0.113\,868$\pm$\$0.032\,088\$0.113 868 ± 0.032 088 0.220 04±0.044 409plus-or-minus0.220040.044409\$0.220\,04$\pm$\$0.044\,409\$0.220 04 ± 0.044 409 0.200 878±0.042 151plus-or-minus0.2008780.042151\$0.200\,878$\pm$\$0.042\,151\$0.200 878 ± 0.042 151 0.243 36±0.070 934plus-or-minus0.243360.070934\$0.243\,36$\pm$\$0.070\,934\$0.243 36 ± 0.070 934 0.137 754±0.044 527plus-or-minus0.1377540.044527\$0.137\,754$\pm$\$0.044\,527\$0.137 754 ± 0.044 527
400 0.226 48±0.051 626plus-or-minus0.226480.051626\$0.226\,48$\pm$\$0.051\,626\$0.226 48 ± 0.051 626 0.104 433±0.028 986plus-or-minus0.1044330.028986\$0.104\,433$\pm$\$0.028\,986\$0.104 433 ± 0.028 986 0.263 64±0.056 862plus-or-minus0.263640.056862\$0.263\,64$\pm$\$0.056\,862\$0.263 64 ± 0.056 862 0.128 294±0.031 032plus-or-minus0.1282940.031032\$0.128\,294$\pm$\$0.031\,032\$0.128 294 ± 0.031 032 0.2826±0.065 494plus-or-minus0.28260.065494\$0.2826$\pm$\$0.065\,494\$0.2826 ± 0.065 494 0.092 855±0.048 178plus-or-minus0.0928550.048178\$0.092\,855$\pm$\$0.048\,178\$0.092 855 ± 0.048 178
500 0.227 36±0.065 321plus-or-minus0.227360.065321\$0.227\,36$\pm$\$0.065\,321\$0.227 36 ± 0.065 321 0.102 41±0.028 423plus-or-minus0.102410.028423\$0.102\,41$\pm$\$0.028\,423\$0.102 41 ± 0.028 423 0.386 96±0.283 25plus-or-minus0.386960.28325\$0.386\,96$\pm$\$0.283\,25\$0.386 96 ± 0.283 25 0.086 264±0.026 684plus-or-minus0.0862640.026684\$0.086\,264$\pm$\$0.026\,684\$0.086 264 ± 0.026 684 0.352 44±0.155 268plus-or-minus0.352440.155268\$0.352\,44$\pm$\$0.155\,268\$0.352 44 ± 0.155 268 0.067 721±0.021 766plus-or-minus0.0677210.021766\$0.067\,721$\pm$\$0.021\,766\$0.067 721 ± 0.021 766
DoorKey 100 0.330 94±0.071 19plus-or-minus0.330940.07119\$0.330\,94$\pm$\$0.071\,19\$0.330 94 ± 0.071 19 0.510 42±0.077 026plus-or-minus0.510420.077026\$0.510\,42$\pm$\$0.077\,026\$0.510 42 ± 0.077 026 0.609 493±0.135 071plus-or-minus0.6094930.135071\$0.609\,493$\pm$\$0.135\,071\$0.609 493 ± 0.135 071 0.325 874±0.070 197plus-or-minus0.3258740.070197\$0.325\,874$\pm$\$0.070\,197\$0.325 874 ± 0.070 197 0.701 847±0.071 458plus-or-minus0.7018470.071458\$0.701\,847$\pm$\$0.071\,458\$0.701 847 ± 0.071 458 0.300 172±0.019 691plus-or-minus0.3001720.019691\$0.300\,172$\pm$\$0.019\,691\$0.300 172 ± 0.019 691
200 0.427 243±0.093 497plus-or-minus0.4272430.093497\$0.427\,243$\pm$\$0.093\,497\$0.427 243 ± 0.093 497 0.493 205±0.027 703plus-or-minus0.4932050.027703\$0.493\,205$\pm$\$0.027\,703\$0.493 205 ± 0.027 703 0.720 069±0.091 289plus-or-minus0.7200690.091289\$0.720\,069$\pm$\$0.091\,289\$0.720 069 ± 0.091 289 0.275 265±0.073 504plus-or-minus0.2752650.073504\$0.275\,265$\pm$\$0.073\,504\$0.275 265 ± 0.073 504 0.846 412±0.037 362plus-or-minus0.8464120.037362\$0.846\,412$\pm$\$0.037\,362\$0.846 412 ± 0.037 362 0.235 007±0.037 117plus-or-minus0.2350070.037117\$0.235\,007$\pm$\$0.037\,117\$0.235 007 ± 0.037 117
300 0.509 762±0.110 659plus-or-minus0.5097620.110659\$0.509\,762$\pm$\$0.110\,659\$0.509 762 ± 0.110 659 0.450 234±0.080 155plus-or-minus0.4502340.080155\$0.450\,234$\pm$\$0.080\,155\$0.450 234 ± 0.080 155 0.757 019±0.122 807plus-or-minus0.7570190.122807\$0.757\,019$\pm$\$0.122\,807\$0.757 019 ± 0.122 807 0.411 938±0.055 965plus-or-minus0.4119380.055965\$0.411\,938$\pm$\$0.055\,965\$0.411 938 ± 0.055 965 0.895 01±0.015 27plus-or-minus0.895010.01527\$0.895\,01$\pm$\$0.015\,27\$0.895 01 ± 0.015 27 0.268 422±0.079 03plus-or-minus0.2684220.07903\$0.268\,422$\pm$\$0.079\,03\$0.268 422 ± 0.079 03
DynamicObstacles 100 0.428 543±0.465 309plus-or-minus0.4285430.465309\$0.428\,543$\pm$\$0.465\,309\$0.428 543 ± 0.465 309 0.044 771±0.037 028plus-or-minus0.0447710.037028\$0.044\,771$\pm$\$0.037\,028\$0.044 771 ± 0.037 028 0.824 501±0.182 298plus-or-minus0.8245010.182298\$0.824\,501$\pm$\$0.182\,298\$0.824 501 ± 0.182 298 0.041 619±0.002 449plus-or-minus0.0416190.002449\$0.041\,619$\pm$\$0.002\,449\$0.041 619 ± 0.002 449 0.335 917±0.460 536plus-or-minus0.3359170.460536\$0.335\,917$\pm$\$0.460\,536\$0.335 917 ± 0.460 536 0.053 889±0.049 941plus-or-minus0.0538890.049941\$0.053\,889$\pm$\$0.049\,941\$0.053 889 ± 0.049 941
200 0.756 247±0.425 772plus-or-minus0.7562470.425772\$0.756\,247$\pm$\$0.425\,772\$0.756 247 ± 0.425 772 0.029 241±0.032 208plus-or-minus0.0292410.032208\$0.029\,241$\pm$\$0.032\,208\$0.029 241 ± 0.032 208 0.986 525±0.020 203plus-or-minus0.9865250.020203\$0.986\,525$\pm$\$0.020\,203\$0.986 525 ± 0.020 203 0.006 841±0.003 298plus-or-minus0.0068410.003298\$0.006\,841$\pm$\$0.003\,298\$0.006 841 ± 0.003 298 0.758 229±0.424 328plus-or-minus0.7582290.424328\$0.758\,229$\pm$\$0.424\,328\$0.758 229 ± 0.424 328 0.030 011±0.019 832plus-or-minus0.0300110.019832\$0.030\,011$\pm$\$0.019\,832\$0.030 011 ± 0.019 832
300 0.779 729±0.435 969plus-or-minus0.7797290.435969\$0.779\,729$\pm$\$0.435\,969\$0.779 729 ± 0.435 969 0.012 641±0.009 381plus-or-minus0.0126410.009381\$0.012\,641$\pm$\$0.009\,381\$0.012 641 ± 0.009 381 0.992 074±0.014 955plus-or-minus0.9920740.014955\$0.992\,074$\pm$\$0.014\,955\$0.992 074 ± 0.014 955 0.000 522±0.000 476plus-or-minus0.0005220.000476\$0.000\,522$\pm$\$0.000\,476\$0.000 522 ± 0.000 476 0.969 484±0.014 053plus-or-minus0.9694840.014053\$0.969\,484$\pm$\$0.014\,053\$0.969 484 ± 0.014 053 0.020 619±0.015 441plus-or-minus0.0206190.015441\$0.020\,619$\pm$\$0.015\,441\$0.020 619 ± 0.015 441
Boxing 1000 0.841 488±0.196 579plus-or-minus0.8414880.196579\$0.841\,488$\pm$\$0.196\,579\$0.841 488 ± 0.196 579 0.346 016±0.1168plus-or-minus0.3460160.1168\$0.346\,016$\pm$\$0.1168\$0.346 016 ± 0.1168 0.931 203±0.053 877plus-or-minus0.9312030.053877\$0.931\,203$\pm$\$0.053\,877\$0.931 203 ± 0.053 877 0.253 524±0.113 155plus-or-minus0.2535240.113155\$0.253\,524$\pm$\$0.113\,155\$0.253 524 ± 0.113 155 0.826 612±0.236 548plus-or-minus0.8266120.236548\$0.826\,612$\pm$\$0.236\,548\$0.826 612 ± 0.236 548 0.252 284±0.046 222plus-or-minus0.2522840.046222\$0.252\,284$\pm$\$0.046\,222\$0.252 284 ± 0.046 222
2000 0.883 905±0.124 857plus-or-minus0.8839050.124857\$0.883\,905$\pm$\$0.124\,857\$0.883 905 ± 0.124 857 0.188 792±0.106 205plus-or-minus0.1887920.106205\$0.188\,792$\pm$\$0.106\,205\$0.188 792 ± 0.106 205 0.933 527±0.057 27plus-or-minus0.9335270.05727\$0.933\,527$\pm$\$0.057\,27\$0.933 527 ± 0.057 27 0.136 189±0.037 203plus-or-minus0.1361890.037203\$0.136\,189$\pm$\$0.037\,203\$0.136 189 ± 0.037 203 0.832 423±0.109 487plus-or-minus0.8324230.109487\$0.832\,423$\pm$\$0.109\,487\$0.832 423 ± 0.109 487 0.300 118±0.182 217plus-or-minus0.3001180.182217\$0.300\,118$\pm$\$0.182\,217\$0.300 118 ± 0.182 217
3000 0.865 427±0.195 112plus-or-minus0.8654270.195112\$0.865\,427$\pm$\$0.195\,112\$0.865 427 ± 0.195 112 0.318 058±0.260 921plus-or-minus0.3180580.260921\$0.318\,058$\pm$\$0.260\,921\$0.318 058 ± 0.260 921 0.825 683±0.143 459plus-or-minus0.8256830.143459\$0.825\,683$\pm$\$0.143\,459\$0.825 683 ± 0.143 459 0.201 04±0.212 373plus-or-minus0.201040.212373\$0.201\,04$\pm$\$0.212\,373\$0.201 04 ± 0.212 373 0.957 467±0.039 969plus-or-minus0.9574670.039969\$0.957\,467$\pm$\$0.039\,969\$0.957 467 ± 0.039 969 0.142 976±0.099 212plus-or-minus0.1429760.099212\$0.142\,976$\pm$\$0.099\,212\$0.142 976 ± 0.099 212
Pong 1000 0.911 429±0.115 539plus-or-minus0.9114290.115539\$0.911\,429$\pm$\$0.115\,539\$0.911 429 ± 0.115 539 0.737 485±0.070 762plus-or-minus0.7374850.070762\$0.737\,485$\pm$\$0.070\,762\$0.737 485 ± 0.070 762 0.836 429±0.205 476plus-or-minus0.8364290.205476\$0.836\,429$\pm$\$0.205\,476\$0.836 429 ± 0.205 476 0.683 161±0.063 495plus-or-minus0.6831610.063495\$0.683\,161$\pm$\$0.063\,495\$0.683 161 ± 0.063 495 0.744 762±0.206 179plus-or-minus0.7447620.206179\$0.744\,762$\pm$\$0.206\,179\$0.744 762 ± 0.206 179 0.734 207±0.051 311plus-or-minus0.7342070.051311\$0.734\,207$\pm$\$0.051\,311\$0.734 207 ± 0.051 311
3000 0.982 381±0.015 622plus-or-minus0.9823810.015622\$0.982\,381$\pm$\$0.015\,622\$0.982 381 ± 0.015 622 0.576 657±0.081 791plus-or-minus0.5766570.081791\$0.576\,657$\pm$\$0.081\,791\$0.576 657 ± 0.081 791 0.982 857±0.005 746plus-or-minus0.9828570.005746\$0.982\,857$\pm$\$0.005\,746\$0.982 857 ± 0.005 746 0.635 608±0.034 63plus-or-minus0.6356080.03463\$0.635\,608$\pm$\$0.034\,63\$0.635 608 ± 0.034 63 0.962 143±0.041 933plus-or-minus0.9621430.041933\$0.962\,143$\pm$\$0.041\,933\$0.962 143 ± 0.041 933 0.599 326±0.037 327plus-or-minus0.5993260.037327\$0.599\,326$\pm$\$0.037\,327\$0.599 326 ± 0.037 327
5000 0.98±0.017 924plus-or-minus0.980.017924\$0.98$\pm$\$0.017\,924\$0.98 ± 0.017 924 0.536 369±0.034 322plus-or-minus0.5363690.034322\$0.536\,369$\pm$\$0.034\,322\$0.536 369 ± 0.034 322 0.942 857±0.092 857plus-or-minus0.9428570.092857\$0.942\,857$\pm$\$0.092\,857\$0.942 857 ± 0.092 857 0.566 152±0.053 357plus-or-minus0.5661520.053357\$0.566\,152$\pm$\$0.053\,357\$0.566 152 ± 0.053 357 0.942 143±0.045 617plus-or-minus0.9421430.045617\$0.942\,143$\pm$\$0.045\,617\$0.942 143 ± 0.045 617 0.604 967±0.043 498plus-or-minus0.6049670.043498\$0.604\,967$\pm$\$0.043\,498\$0.604 967 ± 0.043 498
Table 8: Performance of Sequential-Q, Disagreement-Q, and Random-Q for all environments across different budgets. The reward is reported as the fraction of the reward upper bound. For PixelCartPole, $c$ error is MSE; for the other environments, $c$ error is 1 - accuracy. The $\pm$ [value] part shows the standard deviation. This shows a more complete version of the results in Figure 3.